OpenStudy (anonymous):

is x+6=y^2 a function? How do I figure out if it is? thx.

OpenStudy (nottim):

Well, define function for me.

Directrix (directrix):

x+6=y^2 a function A function has for each x of the domain of the function exactly one y for that x. In x + 6 = y^2, find the y value(s) that go with x when x = 0. If there are two different y values for the same x, then x+6=y^2 is not a function. Graph the equation and see if the graph passes the vertical line test. http://www.wolframalpha.com/input/?i=+x%2B6%3Dy%5E2+

OpenStudy (istim):

No it's not. As directix as had, a function has for each x of the doman of the function exactly one y for that x. Solving for y; y^2 = x + 6 Take the square root, remember when yo utake the square root of a variable, you must account for plus or minus roots. y = plus/minus(x+6) All this means is, for EVERY y, there is a negative answer, and a positive answer. Meaning there are two results for each Y, this means y^2 = x + 6 is NOT a function.